# FlowR: Flowing from Sparse to Dense 3D Reconstructions

### v1

@@ 4.1 Comparison to State-of-the-Art
In Tab. 1, we compare to state-of-the-art sparse-view methods on the DL3DV [29] benchmark. In particular, we compare to InstantSplat [17] and ViewCrafter [64]. For a fair comparison, we use the same prior information (i.e. posed images) and downstream pipelines (i.e. our refined reconstruction) if applicable. While InstantSplat performs competitively to our initial reconstructions in the 12-view setting, its performance stagnates in the 24-view setting, highlighting its limitation to sparse view scenarios. The generated views from ViewCrafter do not improve reconstructions meaningfully in our experiments, as its results with our downstream pipeline lag behind our initial reconstructions. Our full method on the other hand improves 3D reconstruction results by a significant margin across all metrics, both in the 12-view and 24-view settings. The improvement is particularly pronounced in the 12-view setup, and furthermore, the relative gain compared to our initial reconstruction is the highest in perceptual quality (LPIPS). Additionally, we report the results of our flow model applied to the test view poses (FlowR+++++ +) to highlight its effectiveness for refining renderings of unseen views.
@@
Table 2: Dense-view 3D reconstruction on ScanNet+++++ + validation [62]. We use the official training and testing splits. In addition, we report results of our flow-matching model refining the test views (FlowR++), similar to [45] (w/ GAN).
In Tab. 2, we compare to state-of-the-art dense-view methods on the ScanNet+++++ + benchmark. This setting is very challenging, as the initial reconstruction benefits from a large set of source images, and the test view poses are specifically chosen to represent out-of-distribution viewpoints [62]. Despite these challenges, our method improves significantly over our initial reconstruction results, outperforming GANeRF [45]. When applying post-hoc refinement of rendered views (bottom two rows), we also outperform GANeRF by a large margin, improving our initial results dramatically with a most significant improvement in LPIPS. In Fig. 3 and Fig. 4, we show qualitative comparisons on the aforementioned benchmarks. Our method exhibits strikingly fewer artifacts from floaters and poorly reconstructed geometry than the baseline methods.
@@ 4.2 Ablation Studies
Figure 4: Qualitative reconstruction results on ScanNet+++++ + [62]. We show out-of-distribution test view renderings of the baselines and our method in the dense-view setting. This is a particularly challenging setting as the test views are far from the initial camera poses.
@@ 5 Related Work
Generative priors. Generative priors enable improvements in both scene geometry and appearance. While earlier work used normalizing flows [38] and GANs [45], many recent works focus on diffusion models [43, 56, 58, 19, 64, 12, 33]. In particular, many works incorporate diffusion models in the training of radiance fields [43, 56, 58], e.g. via distillation [43, 56]. However, since this requires continuously evaluating the model during training, many recent approaches focus on directly generating additional training views [19, 64, 33]. Among these, many works focus on multi-frame architectures since such models have shown great promise in video generation [4, 5, 9, 60, 34, 42], generating videos with high geometric consistency [28]. However, due to architectural limitations, these works only generate a handful of views at a time. Additionally, they keep the generative process as is and inject prior information into the model only as a conditioning. We instead propose a flow matching process that directly refines initial reconstruction results and design the model architecture such that it still benefits from large-scale text-to-image pre-training.
@@ A.1 Data Details
To collect our large-scale dataset of image pairs, we utilize the large-scale NVS benchmarks DL3DV10K [29] and ScanNet+++++ + [62]. For DL3DV10K, we select $k\in[6,36]$ equally spaced views as the initial sparse training set. We use the 960P resolution images and undistort them before reconstruction. We filter sequences based on scale factor plausibility, i.e. we remove sequences with very large or small scale factors. This makes our calibration pipeline robust to inaccuracies in initial camera parameters or depth estimates. We store the reconstructions in a compressed format following [37]. For ScanNet+++++ +, we first use farthest point sampling to determine a small number of keyframes and subsequently select 25% - 50% of the remaining training set closest to these keyframes. This ensures good spatial coverage while also allowing for target views far from the initial training views, even in a dense view setting. Additionally, we use the out-of-distribution test views for all training sequences. We also undistort the images and resize them to $640\times 960$ resolution.
@@ A.2 Method Details
Initial reconstruction. We check if too few reliable feature tracks can be inferred, and if there are not enough feature tracks, we resort to global pointcloud alignment [27] to ensure sufficiently dense 3D geometry estimates. This is usually only the case for very sparse input view scenarios, i.e. 9 input views or less.
@@
Architecture details. We show an illustration of our multi-view DiT block architecture Fig. 5. As mentioned in  Sec. 3.3, we keep the first self-attention layer, the normalization layers, and the feed-foward layer equivalent to [15], and insert a multi-view attention layer after the first attention layer. This multi-view attention is concluded with a zero linear layer to keep the initialization intact. As mentioned in Sec. 3.2, for each source input view $i$, we condition the model on its camera pose $\mathbf{P}_{i}$ and intrinsics $\mathbf{K}_{i}$, i.e. a pixel $\mathbf{p}^{\prime}$ of image $i$ is represented as a ray $\mathbf{r}=\langle\mathbf{o}\times\mathbf{d},\mathbf{d}\rangle$ with $\mathbf{o}=\mathbf{R}_{j}^{\top}(\mathbf{t}_{i}-\mathbf{t}_{j})$ and $\mathbf{d}=\mathbf{R}_{j}^{\top}\mathbf{R}_{i}\mathbf{K}_{i}^{-1}\mathbf{p}^{\prime}$ where $j$ is the reference view defining common coordinate frame.
@@
Runtime analysis. As mentioned in Sec. 3.1, our initial reconstruction takes an average of 6.5 minutes for pointcloud initialization (6 minutes) and initial 3DGS training (30 seconds). Our multi-view flow model takes approximately 1.5 minutes to generate 200 additional images, processing 45 views at $540\times 960$ resolution (91K tokens) on one H100 GPU at each forward pass. The final reconstruction training takes on average 42.4 minutes, as we use a longer, 30k step schedule and an additional LPIPS loss for the target views.
Limitations. While FlowR makes a significant step towards high-quality, photo-realistic 3D reconstructions from non-exhaustive captures, there remain meaningful directions for future work. For instance, our method relies on heuristics to select the camera views which are used to refine the 3D reconstruction results. In this regard, incorporating uncertainty quantification [23, 20] and active view selection [39] could improve results. Additionally, because our method aims to map incorrect renderings to ground-truth images, its performance depends on the initial 3D reconstruction. In particular, if there are large areas entirely unseen in the source views, our model will not hallucinate new content. Incorporating a proper prior distribution for such cases opens up a promising avenue for future research.
@@ A.3 Additional Experiments
Evaluation details. We use LPIPS with VGG-16 features unless otherwise specified in the benchmark, i.e. we use VGG-16 for all experiments except for the Nerfbusters benchmark which uses AlexNet. Note that for Nerfbusters we resort to the test trajectory for target view selection since it is entirely disconnected from the initial reconstruction and as such it is not possible to do an effective refinement by interpolating along the training trajectory or by sampling poses around it. Finally, we optionally apply naive opacity thresholding, i.e. we define a single minimum opacity value applied to all rendered views to achieve a comparable coverage to our baselines. The intuition behind this is that high opacity along a pixel ray usually correlates with well-defined scene geometry.

### v2

@@ 4.1 Comparison to State-of-the-Art
In Tab. 1, we compare to state-of-the-art sparse-view methods on the DL3DV [30] benchmark. In particular, we compare to InstantSplat [17] and ViewCrafter [67]. For a fair comparison, we use the same prior information (i.e. posed images) and downstream pipelines (i.e. our refined reconstruction) if applicable. While InstantSplat performs competitively to our initial reconstructions in the 12-view setting, its performance stagnates in the 24-view setting, highlighting its limitation to sparse view scenarios. The generated views from ViewCrafter do not improve reconstructions meaningfully in our experiments, as its results with our downstream pipeline lag behind our initial reconstructions. Our full method on the other hand improves 3D reconstruction results by a significant margin across all metrics, both in the 12-view and 24-view settings. The improvement is particularly pronounced in the 12-view setup, and furthermore, the relative gain compared to our initial reconstruction is the highest in perceptual quality (LPIPS). Additionally, we report the results of our flow model applied to the test view poses (FlowR++++) to highlight its effectiveness for refining renderings of unseen views.
@@
Table 2: Dense-view 3D reconstruction on ScanNet++++ validation [65]. We use the official training and testing splits. In addition, we report results of our flow-matching model refining the test views (FlowR++), similar to [47] (w/ GAN).
In Tab. 2, we compare to state-of-the-art dense-view methods on the ScanNet++++ benchmark. This setting is very challenging, as the initial reconstruction benefits from a large set of source images, and the test view poses are specifically chosen to represent out-of-distribution viewpoints [65]. Despite these challenges, our method improves significantly over our initial reconstruction results, outperforming GANeRF [47]. When applying post-hoc refinement of rendered views (bottom two rows), we also outperform GANeRF by a large margin, improving our initial results dramatically with a most significant improvement in LPIPS. In Fig. 3 and Fig. 4, we show qualitative comparisons on the aforementioned benchmarks. Our method exhibits strikingly fewer artifacts from floaters and poorly reconstructed geometry than the baseline methods.
@@ 4.2 Ablation Studies
Figure 4: Qualitative reconstruction results on ScanNet++++ [65]. We show out-of-distribution test view renderings of the baselines and our method in the dense-view setting. This is a particularly challenging setting as the test views are far from the initial camera poses.
@@ 5 Related Work
Generative priors. Generative priors enable improvements in both scene geometry and appearance. While earlier work used normalizing flows [39] and GANs [47], many recent works focus on diffusion models [45, 59, 61, 19, 67, 12, 34]. In particular, many works incorporate diffusion models in the training of radiance fields [45, 59, 61], e.g. via distillation [45, 59]. However, since this requires continuously evaluating the model during training, many recent approaches focus on directly generating additional training views [19, 67, 34, 41]. Among these, many works focus on multi-frame architectures since such models have shown great promise in video generation [4, 5, 9, 63, 35, 44], generating videos with high geometric consistency [29]. However, due to architectural limitations, these works only generate a handful of views at a time. Additionally, they keep the generative process as is and inject prior information into the model only as conditioning. We instead propose a flow matching process that directly refines initial reconstruction results and design the model architecture such that it still benefits from large-scale text-to-image pre-training.
@@ A.1 Data Details
To collect our large-scale dataset of image pairs, we utilize DL3DV10K [30] and ScanNet++++ [65] benchmarks. For DL3DV10K, we select $k\in[6,36]$ equally spaced views as the initial sparse training set. We use the 960P resolution images and undistort them before reconstruction. We filter sequences based on scale factor plausibility, i.e. we remove sequences with very large or small scale factors. This makes our calibration pipeline robust to inaccuracies in initial camera parameters or depth estimates. We store the reconstructions in a compressed format following [38]. For ScanNet++++, we first use farthest point sampling to determine a small number of keyframes and subsequently select 25% - 50% of the remaining training set closest to these keyframes. This ensures good spatial coverage while also promoting target views far from the initial training views, even in a dense view setting. Additionally, we use the out-of-distribution test views of the training sequences. We undistort the images and resize them to $640\times 960$. The resulting dataset covers a wide variety of input rendering quality for training, with initial PSNR values ranging from 5 to more than 30.
@@ A.2 Method Details
Initial reconstruction. Note that before running point triangulation, we check if there are sufficient reliable feature tracks inferred from the MASt3R matches, and if there are not enough feature tracks, we resort to global pointcloud alignment [28] to ensure sufficiently dense 3D geometry estimates. This is usually only the case for very sparse input view scenarios, i.e. 9 input views or less.
@@
Num. views GPU mem. (GB) PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
72 42.3 24.15 0.808 0.180
12 17.0 24.29 0.801 0.175
Table 7: Memory consumption vs. performance across input view counts. We show our model is effective both at high and low number of views on DL3DV140 at 512px width (cf. Tab. 5).
@@
Architecture details. The overall architecture of our model follows [15]. To compress the input images into latents, we use a VAE [48] with latent dimension of 16. The latents are patchified with a patch size of 2 and fed to our latent flow matching model, which consists of 24 DiT blocks [42]. Each block has 24 attention heads with dimension 64, and the feedforward network has the same hidden dimension of 1536. We show an illustration of our multi-view DiT block architecture Fig. 5. As mentioned in  Sec. 3.3, we keep the first self-attention layer, the normalization layers, and the feed-foward layer equivalent to [15], and insert a multi-view attention layer after the first attention layer. This multi-view attention is concluded with a zero linear layer to keep the initialization intact. As mentioned in Sec. 3.2, for each source input view $i$, we condition the model on its camera pose $\mathbf{P}_{i}$ and intrinsics $\mathbf{K}_{i}$, i.e. a pixel $\mathbf{p}^{\prime}$ of image $i$ is represented as a ray $\mathbf{r}=\langle\mathbf{o}\times\mathbf{d},\mathbf{d}\rangle$ with $\mathbf{o}=\mathbf{R}_{j}^{\top}(\mathbf{t}_{i}-\mathbf{t}_{j})$ and $\mathbf{d}=\mathbf{R}_{j}^{\top}\mathbf{R}_{i}\mathbf{K}_{i}^{-1}\mathbf{p}^{\prime}$ where $j$ is the reference view defining common coordinate frame.
Runtime and memory analysis. As mentioned in Sec. 3.1, our initial reconstruction takes an average of 6.5 minutes for pointcloud initialization (6 minutes) and initial 3DGS training (30 seconds). Our multi-view flow model takes approximately 1.5 minutes to generate 200 additional images, processing 45 views at $540\times 960$ resolution (91K tokens) on one H100 GPU at each forward pass. The final reconstruction training takes on average 42.4 minutes, as we use a longer, 30k step schedule and an additional LPIPS loss for the target views. In Tab. 7, we report the memory usage of our model and show that it can run both on high-end and memory-constrained consumer grade GPUs by adjusting the number of input views that are processed in a single forward pass. Our model reaches comparable performance on our DL3DV140 benchmark with 72 and 12 views, respectively.
Limitations. While FlowR makes a significant step towards high-quality, photo-realistic 3D reconstructions from non-exhaustive captures, there remain meaningful directions for future work. For instance, our method relies on heuristics to select the camera views that are used to refine the 3D reconstruction results. In this regard, incorporating uncertainty quantification [24, 20] and active view selection [40] could improve results. Additionally, because our method aims to map incorrect renderings to ground-truth images, its performance depends on the initial 3D reconstruction. In particular, if there are large areas entirely unseen in the source views, our model will not hallucinate new content. Incorporating a suitable prior for such cases opens up a promising avenue for future research. Finally, we focus on static scenes, as 3DGS exhibits major artifacts with dynamic objects. Hence, another meaningful direction for future work is to extend our method to dynamic scenes using a dynamic 3DGS method, which requires adapting the training data to reflect the presence of dynamic objects in both the source and target distributions.
@@ A.3 Additional Experiments
Evaluation details. We use LPIPS with VGG-16 features unless otherwise specified in the benchmark, i.e. we use VGG-16 for all experiments except for the Nerfbusters benchmark which uses AlexNet. Note that for Nerfbusters we resort to the test trajectory for target view selection since the test views are entirely disconnected from the initial reconstruction and as such it is not possible to effectively refine the reconstructed scene by interpolating along the training trajectory or by sampling poses around it. Finally, we optionally apply naive opacity thresholding, i.e. we define a single minimum opacity value applied to all rendered views to achieve a comparable coverage to our baselines. The intuition behind this is that high opacity along a pixel ray usually correlates with well-defined scene geometry.